OpenStudy (countrygirl1431):

Multiply or divide as indicated. r 2 • r • r 5

OpenStudy (countrygirl1431):

@Nnesha this would be r^10 right of is it somthing different

OpenStudy (anonymous):

the anser is R^8

Nnesha (nnesha):

no how is it 10?

OpenStudy (countrygirl1431):

2*5=10

Nnesha (nnesha):

just add the exponent

Nnesha (nnesha):

no no!

Nnesha (nnesha):

when you `multiply` same bases you should `add` their exponents

Nnesha (nnesha):

\[\huge\rm x^m \times x^n \times x^z=x^{m+n+z}\] just like you did on the previous post

OpenStudy (countrygirl1431):

oh ok i got it now thanks

Nnesha (nnesha):

np :=(